Calculate the molarity of a concentrated H2SO4 solution of density 1.84 g/ml and
containing 98.0 % H2SO4 solute by mass
1
Expert's answer
2019-06-18T05:47:07-0400

"m(H2SO4)=0.98*1.84g\/mol*1000mL=1803.2g"

"n(H2SO4)=\dfrac{1803.2g}{98g/mol}=18.4moles"

"c(H2SO4)=\dfrac{18.4moles}{1L}=18.4mol/L"